Use implicit differentiation to find the following. (Round answers to four decimal places as needed. If only th (xy)2 + xy -Use implicit differentiation to find the following. (Round answers to four decimal place In(x - y) + 2 = 2x2, x = 0 (a) the eParamount Electronics has an annual protit given by P= -100,000 + 5,0009 - 0.2592 dollars, where is the number of laptop compThe percentage y (of total personal consumption) an individual spends on food is approximately y = 33x0.63 percentage points이 Find the equation of the tangent line to the graph of the given function at the point with the indicated x-coordinate. F(x)Find the equation of the tangent line to the graph of the given function at the point with the indicated x-coordinate. F(x) =
Answer #1

☺ @ Gwen (ay)° +xy - x = 3, (3, 0) = x@y? + xy -X =3 differentialing wort. x, we get 2x? y dy + 2xy2 + x dy +y +=0 (184 +9) d- 4x + 484 + at to, An(.y) = 822-1) m-y = 62(x+1) 3 ya x-e? 5 ye 0-8761) = 42 so 3. Equation of tangent is 9 + 1 = 1 (2-0) ar( | ROR xt 3 | SI-2 Gven fra) = X ,=2 | so y = ft) - 2+ 그 C and f(x) = 1 - = 1 - 4 so the equation of fangent line is, y - k
Answer #2

ac y - x = 3 0 (0) (2+ -4-5, (-3,0) (0) (ayyºr ay - X=3 > x*y? + xy - x = 3 O The expression of the slope of the tangent linedx > day 9 + day + 20y + y-t=0 > (22yt a ) + dayty-two => (da?y + x) dx = -dxy-y #1 (6-) line at the The equation of the tm=) at(-1,0) m = V +1 + 2 lat (-3,0) -3 x 3)xo* - 0 +1 2(-3)*x0 + 6-3) 0-0 + 1 (3) - Scanned with CamScanner